5. The correct designation of the bonds which links the CH to the CH, in CH=CH-CH, is: (a) sp-sp sigma (b) sp-sp sigma (c)
Answer #1

5. hy briclization of 6-2 = No. of bond be Ayo? A No. of lone fair hybridization of c-3 = 3+0 = sp2 = No. of o-bond No. of lHybridisation calculation : number of sigma bonds + number of lone pairs (x)

X: 2(sp) , 3(sp2), 4(sp3)

*Note:

1. Single bond has only 1 sigma bond

2. Double bond has 1 sigma bond and 1 pi- bond.

3. Triple bond has 1 sigma bond and 2 pi- bond.